OpenStudy (anonymous):

9.3 Resolve the force vector F = (990 N)i - (330 N) j + (660 N)k into two components, one parallel to vector a = i + 3k , the other perpendicular to a .

OpenStudy (ash2326):

@Arnab09 would you please explain how you got it?

OpenStudy (anonymous):

yeah, obviously..

OpenStudy (ash2326):

Your perpendicular component's dot product with a is not zero!!!

OpenStudy (anonymous):

lemme check again

OpenStudy (ash2326):

No worries:)

OpenStudy (anonymous):

oh sorry, my method was wrong.. :( trying it afresh

OpenStudy (anonymous):

yeah, answer is coming: 297Ni+891Nj (parallel) 693Ni-330Nj-231Nk (perpendicular)

OpenStudy (anonymous):

@ash2326 , check it :)

OpenStudy (ash2326):

Isn't the sum of these two vectors be F?

OpenStudy (anonymous):

yes, of course..

OpenStudy (ash2326):

F=990i-330j+660k a=i+3k Let's find the angle between F and a let A be unit vector parallel to a \[A= (a)/|a|=\frac{i+3k}{\sqrt {10}}\] let Pa be F's component parallel to a \[Pa= (F.A)A\] \[Pa=((990i-330j+660k).\frac{i+3k}{\sqrt {10}})\frac{i+3k}{\sqrt {10}}\] \[Pa=(990+1980)\frac{i+3k}{ 10}\] \[Pa=297i+891j\] Perpendicular component Pe \[Pe=F-Pa=990i-330j+660k-(297i+891k)\] \[Pe=693i-330j-231k\] So parallel Component=297i+891j Perpendicular Component=693i-330j-231k

OpenStudy (anonymous):

@ash2326 , it can be solved in easier way, i guess

OpenStudy (ash2326):

@Arnab09 great work but next time explain your answer. Just giving answers won't help the user who has asked the question.

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

i solved it this way, say, the perpendicular vector is xi+yj+zk so, dot product of this vector and a will be zero so, after dot product, we get a condition: x+3z=0 also, the other vector is (990-x)i-330j+(660-z)k but this is parallel to a so, we get another condition: 3(990-x)=660-z solving these two, we can get x and z

OpenStudy (anonymous):

also, the parallel vector should not have any j component, as it is parallel to a. so, y=-330